y^2/3 - 10y^1/3 + 9 = 0. I need to solve using quadratic systems.
Found 2 solutions by rothauserc, Alan3354:
Answer by rothauserc(4718)

y^2/3 - 10y^1/3 + 9 = 0
factor the above equation
(y^(1/3) - 9) * (y^(1/3) - 1) = 0
then
y^(1/3) = 9
y = 2.080083823
also
y^(1/3) = 1
y = 1

Answer by Alan3354(69443)

y^2/3 - 10y^1/3 + 9 = 0. I need to solve using quadratic systems.
-----------
The other tutor factored, but that should give you a hint.
---
Sub x for y^(1/3)
--> x^2 - 10x + 9 = 0
etc